OpenStudy (anonymous):

Finding The Slope of A line: A(2,6) D(-3,-3)

OpenStudy (anonymous):

change in y/change in x. Or -3-6/-3-2 = -9/-5 = 9/5.

OpenStudy (anonymous):

finding the coordinate when the slope is given: 1=y-3/5-2